OpenStudy (anonymous):

Twenty-five members of the eighth grade class at Park Center Middle School are going to a museum and then to lunch. Each student must pay an entrance fee to the museum and $7.25 for lunch. The total cost for the trip is $443.75. What is the entrance fee for one student? the answer is 17.46 or 17.75

OpenStudy (anonymous):

@lilg132: Please don't just provide answers. From the code of conduct: "Don't provide someone with just the answer - explain the process, and help guide them through understanding the problem."

OpenStudy (anonymous):

The unknown value is the entrance fee, so let's call it F. Each student pays the entrance fee plus $7.25 for lunch, or: \[F + 7.25\] There are 25 students, so the total paid is: \[25(F + 7.25)\] And we know that the total paid is $443.75. So here we go: \[25(F + 7.25) = 443.75\]\[25F + 25\times7.25 = 443.75\]\[25F + 181.25 = 443.75\]\[25F + 181.25 - 181.25 = 443.75 - 181.25\]\[25F = 262.5\]\[F = \frac{ 262.5 }{ 25 }\]\[F = 10.5\] So the entrance fee is actually $10.50. We can make sure if we want (always a good idea) by plugging the result back into our equation: \[25(F + 7.25)\]\[25(10.50 + 7.25)\]\[25(17.75)\]\[443.75\] So we know that $10.50 is the correct answer!

OpenStudy (anonymous):

thx and um whoever the top one is...da flutter spaming much!